OpenStudy (samanthalaclair):

How do you simplify x5/9 Times x1/3? Help!

OpenStudy (samanthalaclair):

also i need help simplifying 4 over 2-5i

OpenStudy (karleighxxvictoria):

its 2-5i

OpenStudy (samanthalaclair):

thats what i said but my teacher claimed there was more to it -_- , thank god I'm not the only one!

OpenStudy (karleighxxvictoria):

lol sometimes teachers are VERY VERY wrong!

OpenStudy (samanthalaclair):

do you know the answer to the first question i asked?

OpenStudy (karleighxxvictoria):

no sorry

OpenStudy (mathstudent55):

By "x5/9 Times x1/3" do you mean \(\Large x^{\frac{5} {9}} \times x^{\frac{1} {3}} \) ? If so, when you multiply powers with the same base, add the exponents. Keep in mind that 5/9 and 1/3 have different denominators, so you need a common denominator. In this case use 9 as the LCD. For the second problem, you do it depending on what the problem is. You wrote "4 over 2-5i". Do you mean \(\dfrac{4}{2} - 5i\) or do you mean \(\dfrac{4}{2 - 5i} \) ? If you meant \(\dfrac{4}{2} - 5i\), then the answer is \(2 - 5i\). If your teacher told you it's wrong and it's more complicated than that, then you probably mean the second possibility: \(\dfrac{4}{2 - 5i} \) Now you need to use a complex conjugate to simplify the fraction. The complex conjugate of 2 - 5i is 2 + 5i, so multiply the numerator and denominator by 5 + 2i.

OpenStudy (studybuddy26):

How is any of this high school math ? I took all this stuff in middle school. Also including Geometry 1.